84 students who took the achievement test last Oct. 2020 If the mean is 88, s = 2.50. how many students scored above the mean?



1
Expert's answer
2021-11-17T15:21:55-0500

"P(X>88) = 1-P(X<88) \\
= 1 -P(Z< \frac{88-88}{2.5}) \\
= 1 - P(Z<0) \\
= 1 -0.5 \\
= 0.5 \\
n=84 \\
Expected = np = 84 \times 0.5 = 42"